The present value of $13,000 received seven years from today, assuming an interest rate of 5% compounded annually is:

A) 9,238.85

B) 17,550.00

C) 18,292.30

D) 19,500.00
Answer #1

Answer is A) 9,238.85

Using compound interest formula,

A = P(1+i)n , Where A=Future value, P=present value, i = Interest rate & n = No. of years

$13,000 = P(1+0.05)7

13000 = P(1.05)7

P = 13000/1.4071

P   = 9,238.857